OpenStudy (anonymous):

2x/4x+2 * 14x+7/6 a. 7x/12 b. 7/6 c. x/6 d. 7x/6

OpenStudy (anonymous):

Do you mean \[\frac{ 2x }{ 4x + 2 } \times 14x + \frac{ 7 }{ 6 }\] ??

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{ 2x }{ 4x +2 } \times \frac{ 4x + 7 }{ 6 }\] \[\frac{ 2x(4x+7) }{ (4x+2)6 }\] \[\frac{ 8x ^{2} + 14x }{ 24x + 12 }\]